# Tame deformations of highly singular function germs

### v2

@@ 6. Examples
By the following four examples we show here the variety of situations that may occur: the Jacobian criterion of Theorem 5.6 is satisfied in 6.2 but not in 6.1; the condition (3) is satisfied in 6.1 but not in 6.3; and finally, the tameness condition holds in 6.3 but not in 6.4.
@@
Let $F:({\mathbb{K}}^{2}\times{\mathbb{K}},0)\to({\mathbb{K}},0)$, $F(x,y,t)=y^{2}(x^{2}-(y-t)^{2})$. This is a deformation of a line singularity, as considered in [Si1] by Siersma in the complex setting. The function germ $F_{t}$ has the line $\{y=0\}\subset{\mathbb{K}}^{2}$ as singular locus, for any parameter $t$ close enough to 0.
Nevertheless ${\mathrm{Sing\hskip 2.0pt}}F$ consists of two components: one is the surface $\{y=0\}\times{\mathbb{K}}$, and the other is the line $\{x=0,y=t\}$. We have the equality ${\mathrm{Sing\hskip 2.0pt}}F={\mathrm{Sing\hskip 2.0pt}}\widetilde{F}$, thus condition (5) is satisfied. The discriminant $\Delta:=\widetilde{F}({\mathrm{Sing\hskip 2.0pt}}\widetilde{F})$ is the axis $\{0\}\times{\mathbb{K}}$.
By simple computations we see that condition (3) is satisfied, whereas condition (16) is not satisfied precisely at the origin, and therefore the hypothesis of the Jacobian criterion Theorem 5.6 does not hold either. Actually such a deformation of a line singularity with singular locus $L$ is called “admissible” in [Si1], [Si2] (and in more other papers), and has the property that the line $L\subset{\mathrm{Sing\hskip 2.0pt}}F_{t}$ preserves its generic transversal type (which is $A_{1}$ in the above example), while from the origin spring, when $t\not=0$, finitely many special points (i.e. of a different transversal type) inside $L$, as well as finitely many singular points outside $L$.
In the complex setting, it turns out by a proof similar to Lê-Saito’s in [LS], that the $t$-constancy of the generic transversal Milnor number implies condition (3), thus the “admissible” deformations are in fact tame, by our Theorem 4.3.
@@
Inspired by Teissier’s examples in [Te4], let us consider the deformation $F(x,t)=z_{1}^{5}+z_{2}^{5}+z_{1}^{6}z_{2}^{6}z_{3}^{2}+tz_{1}^{3}z_{2}^{3}$ of the polynomial $F_{0}=z_{1}^{5}+z_{2}^{5}+z_{1}^{6}z_{2}^{6}z_{3}^{2}$ with ${\mathrm{Sing\hskip 2.0pt}}F_{0}=\{z_{3}=0\}$, either over ${\mathbb{C}}$ or over ${\mathbb{R}}$. Note that the inclusion $(\partial F_{t})\subset(\partial F_{0})$ doesn’t hold since $z_{1}^{3}z_{2}^{2}\in\overline{(\partial F_{0})}\setminus{(\partial F_{0})}$ and $z_{1}^{2}z_{3}^{2}\in\overline{(\partial F_{0})}\setminus{(\partial F_{0})}$.
@@
In the following deformation the singular locus of $F_{0}$ splits into several pieces “at the boundary”.
@@
The singular locus ${\mathrm{Sing\hskip 2.0pt}}F_{0}$, consisting of the two axes $\{x=y=0\}\cup\{x=z=0\}$, splits indeed, for $t\not=0$, into several curves:
${\mathrm{Sing\hskip 2.0pt}}F_{t}=\{x=y=0\}\cup\{x=\frac{2}{3}t,\ y=0\}\cup\{x=t,\ y^{2}z+x^{2}=0\}.$
One may establish (by straightforward computations) that $F$ satisfies condition (5), does not satisfy condition (3), but that $F$ is tame, cf Definition 3.5. Indeed, the Milnor set:
@@
Our Theorem 3.7 then tells that this deformation is trivial at the boundary, and moreover that $\widetilde{F}$ has a Milnor-Hamm tube fibration.
Let us point out that in [Hof, Example 9.2] the author considers the same $F_{0}$ in the complex analytic setting, within a 2-parameter deformation, and shows that his criterion [Hof, Theorem 1.2] applies; then, by using this deformation, he computes (via the method of [ST2]) the homology of the Milnor fibre of $F_{0}$.
@@
We have $\partial F=(-3x^{2}+2txz,2y,tx^{2};\ x^{2}z)$, thus ${\mathrm{Sing\hskip 2.0pt}}F={\mathrm{Sing\hskip 2.0pt}}\widetilde{F}$, and in particular our condition (5) is satisfied. Note that ${\mathrm{Sing\hskip 2.0pt}}F_{t}$ is constant whereas the Jacobian ideal $\partial F_{t}$ is not constant for $t$ in some neighbourhood of 0.
The computation of the Milnor set over ${\mathbb{R}}$ is easier; we get that $M(\widetilde{F})$ contains a surface in the 3-space $\{y=0\}\subset{\mathbb{R}}^{3}\times{\mathbb{R}}$ of coordinates $x,z,t$, and the intersection of $M(\widetilde{F})$ with the slice $t=0$ contains ${\mathrm{Sing\hskip 2.0pt}}F_{0}=\{(0,0,z,0)\}$. This tells that $\widetilde{F}$ is not $\rho$-regular, which implies that the $\partial$-Thom regularity fails [^4] There are examples of maps which fail to be Thom regular but they are however $\rho$-regular, cf [ACT1]. It remains here the question if this can happen in case of deformations. along the $z$-axis. As we have remarked at the end of the proof of Theorem 3.7, the $\rho$-regularity is equivalent to the fact that the tube map (6) is a proper stratified submersion.
Over ${\mathbb{R}}$, since this example is not $\rho$-regular (i.e. not tame), it cannot satisfy any other criteria implying topological triviality at the boundary. In the complex setting, in [Hof] the author computes explicitly the fibres of the fibration (2) and finds that actually they have different Euler characteristics, confirming the topological non-triviality at the boundary.

### v3

@@ 6. Examples
By the following four examples we show here the variety of situations that may occur: the Jacobian criterion of Theorem 5.6 is satisfied in 6.2 but not in 6.1; the condition (3) is satisfied in 6.1 but not in 6.3; and finally, the tameness condition holds in 6.3 but not in 6.4. We remind here too that, in all our notations, the maps and the sets defined by them are regarded as germs at the origin.
@@
Let $F:({\mathbb{K}}^{2}\times{\mathbb{K}},0)\to({\mathbb{K}},0)$, $F(x,y,t)=y^{2}(x^{2}-(y-t)^{2})$. This is a deformation of a line singularity, in the terminology used by Siersma in [Si1] in the complex setting only. For any parameter $t$ close enough to 0, the function germ $F_{t}$ has indeed as singular locus ${\mathrm{Sing\hskip 2.0pt}}F_{t}:=\{y=0\}\subset{\mathbb{K}}^{2}$, hence the same line.
In our setting, we have the following set germs in $({\mathbb{K}}^{3},0)$: ${\mathrm{Sing\hskip 2.0pt}}F=\{x=0,y=t\}\cup\{y=0\}$, and ${\mathrm{Sing\hskip 2.0pt}}\widetilde{F}=\{x=0,y=t\}\cup\{x=0,2y=t\}\cup\{y=0\}$.
Let us notice the strict inclusions: $\bigcup_{t}{\mathrm{Sing\hskip 2.0pt}}F_{t}=\{y=0\}\subsetneq{\mathrm{Sing\hskip 2.0pt}}F\subsetneq{\mathrm{Sing\hskip 2.0pt}}\widetilde{F}$.
The equality $({\mathrm{Sing\hskip 2.0pt}}F)_{|t=0}=({\mathrm{Sing\hskip 2.0pt}}\widetilde{F})_{|t=0}$ however holds, thus the preliminary condition (6) is satisfied. The discriminant $\Delta:=\widetilde{F}({\mathrm{Sing\hskip 2.0pt}}\widetilde{F})$ is the union of the germ at 0 of the axis $\{0\}\times{\mathbb{K}}$ with the germ of the curve in ${\mathbb{K}}^{2}$ parametrized as $\{(-t^{4}/16,t)\}$.
By simple computations we see that condition (3) is satisfied, thus Theorem 1.3 holds, whereas condition (17) is not satisfied precisely at the origin, and therefore the hypothesis of the Jacobian criterion Theorem 5.6 does not hold either. In the complex setting, such a deformation of a line singularity with singular locus $L$ was called “admissible” in [Si1], [Si2] and in more other papers, and has the property that the line $L={\mathrm{Sing\hskip 2.0pt}}F_{t}$ preserves its generic transversal type (which is $A_{1}$ in the above example), while from the origin may spring, when $t\not=0$, finitely many special points along $L$ (i.e. with a different transversal type), as well as finitely many singular points outside $L$.
In the complex setting, it turns out by a proof similar to Lê-Saito’s in [LS] applied at some generic point $p\in L\setminus\{0\}$, that the $t$-constancy of the generic transversal Milnor number implies condition (3), thus the “admissible” deformations are in fact tame, by our Theorem 4.3.
@@
Let us consider the deformation $F(x,t)=z_{1}^{5}+z_{2}^{5}+z_{1}^{6}z_{2}^{6}z_{3}^{2}+tz_{1}^{3}z_{2}^{3}$ of the polynomial $F_{0}=z_{1}^{5}+z_{2}^{5}+z_{1}^{6}z_{2}^{6}z_{3}^{2}$ with ${\mathrm{Sing\hskip 2.0pt}}F_{0}=\{z_{1}=z_{2}=0\}$, either over ${\mathbb{C}}$ or over ${\mathbb{R}}$. Note that the inclusion $(\partial F_{t})\subset(\partial F_{0})$ doesn’t hold since $z_{1}^{3}z_{2}^{2}\in\overline{(\partial F_{0})}\setminus{(\partial F_{0})}$ and $z_{1}^{2}z_{3}^{2}\in\overline{(\partial F_{0})}\setminus{(\partial F_{0})}$.
@@
The singular locus ${\mathrm{Sing\hskip 2.0pt}}F_{0}$ consists of the germs of the two axes $\{x=y=0\}\cup\{x=z=0\}$, the first of multiplicity 4 and the second of multiplicity 1. For $t\not=0$, the singular locus ${\mathrm{Sing\hskip 2.0pt}}F_{t}$ is the germ at $0\in{\mathbb{R}}^{3}$ of $\{x=y=0\}$ with multiplicity 1. There are two more curves in ${\mathbb{R}}^{3}\times\{t\}$ which emerge from ${\mathrm{Sing\hskip 2.0pt}}F_{0}$:
$\bigl{\{}x=\frac{2}{3}t,\ y=0\bigr{\}}\cup\bigl{\{}x=t,\ x^{2}+y^{2}z=0\bigr{\}}$
that are no more germs at $0\in{\mathbb{R}}^{3}$. Their union over $t$, together with $\cup_{t}{\mathrm{Sing\hskip 2.0pt}}F_{t}$, constitute the singular set germ ${\mathrm{Sing\hskip 2.0pt}}\widetilde{F}$.
By straightforward computations, one establishes that $F$ satisfies condition (6), that $F$ does not satisfy condition (3), but that $F$ is tame, cf Definition 3.5. Indeed, the Milnor set:
@@
Our Theorem 3.7 then tells that this is a deformation with fibre constancy, and moreover that $\widetilde{F}$ has a Milnor-Hamm tube fibration.
Let us point out that in [Hof, Example 9.2] the author considers the same $F_{0}$ in the complex analytic setting (moreover within a 2-parameter deformation) and shows that his algebraic criterion [Hof, Theorem 1.2] applies to conclude that this is a deformation is “admissible” in the terminology of [ST2]. Through this deformation, Hof can then compute the homology of the Milnor fibre of $F_{0}$ by using the special method developed [ST2] in the complex setting.
@@
We have $\partial F=(-3x^{2}+2txz,2y,tx^{2};\ x^{2}z)$, thus ${\mathrm{Sing\hskip 2.0pt}}F={\mathrm{Sing\hskip 2.0pt}}\widetilde{F}$, and in particular our condition (6) is satisfied. Note that the Jacobian ideal $\partial F_{t}$ is not constant for $t$ in some neighbourhood of 0, despite the fact that the set germ ${\mathrm{Sing\hskip 2.0pt}}F_{t}$ is constant.
The computation of the Milnor set over ${\mathbb{R}}$ is easier here; we get that $M(\widetilde{F})$ contains a surface germ in the 3-space $\{y=0\}\subset{\mathbb{R}}^{3}\times{\mathbb{R}}$ of coordinates $x,z,t$, and the intersection of $M(\widetilde{F})$ with the slice $t=0$ contains ${\mathrm{Sing\hskip 2.0pt}}F_{0}=\{(0,0,z,0)\}$. This tells that $\widetilde{F}$ is not $\rho$-regular, which implies that the $\partial$-Thom regularity fails [^4] There are examples of maps which fail to be Thom regular but they are however $\rho$-regular, cf [ACT1]. It remains here the question if this can happen in case of deformations. along the $z$-axis. As we have remarked at the end of the proof of Theorem 3.7, the $\rho$-regularity is equivalent to the fact that the tube map (7) is a proper stratified submersion.
Since this example over ${\mathbb{R}}$ is not $\rho$-regular (i.e. not tame, cf Definition 3.5), it cannot satisfy any other criteria implying deformation with fibre constancy. In the complex setting, Hof remarks in [Hof] that this example does not satisfy his criteria; then he computes explicitly the fibres of the fibration (2) and finds that the homotopy type varies (actually one has $S^{2}$ for $t\not=0$ and $S^{1}\vee S^{1}$ for $t=0$), confirming that the deformation does not have fibre constancy.